Based on the thermodynamic properties provided for water, determine the amount of energy needed for 227 g of water to go from 71.5 °C to 181 °C.
Answer #1

We know that,

Specific heat capacity of water (liquid) = c1 = 4.184 J/°C g

Specific heat capacity of water (vapor) = c2 = 1.996 J/°C g

Enthalpy of vaporization (Hv) = 2260 J/g

Mass of water (m) = 227 g

We can consider below scenarios for overall energy calculation,

1) Temperature increase from 71.5 to 100°C

Q1 = m * c1 * (100 - 71.5) = 227*4.184*28.5 = 27068.4 J

2) phase change from liquid to gas

Q2 = m * Hv = 227 * 2260 = 513020 J

3) Temperature increase from 100 to 181°C

Q3 = m * c2 * (181 - 100) = 227 * 1.996 * 81 = 36700.45 J

Total energy (Q) = Q1 + Q2 + Q3 = 576788.852 J ... Answer